OpenStudy (anonymous):

lim x - 3 x->2+ ---------- x^2 - 4

OpenStudy (anonymous):

what is lim x..>2+????

OpenStudy (anonymous):

lol ,i know but i'am asking about + ??

OpenStudy (anonymous):

Jatin I know :) ,but check out the question again is it written x..>2 ,or , x....>2+

OpenStudy (anonymous):

x is approaching to 2 from Right and we have to find out the Right Hand Limit i'm so idiot in typing. i wil be mad :( :(

OpenStudy (anonymous):

just one side limit we have to find out ...

OpenStudy (anonymous):

OMG *_* ,iknow what is limits ,but i'am asking about the + sign !!!!!

OpenStudy (anonymous):

huhuhuhuhu. u know i'm weeping bcoz i make so much mistake.... ... :( :( :( ... in my mathematics book , the question is find out one sided limit only .......... in this question we have to find out Right hand limit.... where x---> 2 from Right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@eyad. u left me .. :) :)hahahahaah

OpenStudy (anonymous):

hahahahahah ,lol i left my mistake i pressed back by mistake hahahahahahahhaa

OpenStudy (anonymous):

no no .. u r in here .. i thought u left me .. i mean my question hahaahahahah... it's alright :) :)

OpenStudy (anonymous):

@Eyad hope u r going to solve my problem.... :) .. the answer in my buk is -\[\infty\]

OpenStudy (anonymous):

the answer is - infinity

OpenStudy (anonymous):

ok ,i'am solving it lol :)

OpenStudy (anonymous):

ohhh... sure plz carry on ..... thanks alot :)

OpenStudy (anonymous):

first of all u will subst. the x with 2 ,so if u get a real number so its k ,if not u will factorize ,anyway factorizing will be useless ,so when u subist. x with 2 ,u will get 5/0 where = infinity

OpenStudy (anonymous):

yah.. further plz .. i got everything till now u explained to me .. further plz

OpenStudy (anonymous):

i mean any answer be a/0=infinity or undifined

OpenStudy (anonymous):

ummm ,wait i got a paper will help u in limits kinds ...

OpenStudy (anonymous):

@Eyad i like the way u r attempting my question :) :) :)

OpenStudy (anonymous):

look , i captured a paper from a book in limits and i will upload to the pc now ok :)

OpenStudy (anonymous):

thanks thanks thanks thanks .. u r angel to me :) :)

OpenStudy (anonymous):

Here is it ,i hope it helps

OpenStudy (anonymous):

look at the first point ..

OpenStudy (anonymous):

@Eyad .. huhuhuhuhuhuhuhuh.. :( :( :( ..............where is my solution .... :( :( .. i was so happy to know that i will get my solution on paper but there it is not ..... what is ur name .. i wanna be ur friend .. u r so nice even though i didn't get my answer :) :) will u be my friend plz :)

OpenStudy (anonymous):

@Eyad u know that paper helped me a lot really .. :) :)

OpenStudy (anonymous):

@Eyad .... u r mindBlowing... that is really helpful .. there are so many useful big things on just that small paper ... thanks frienddddddddd .. :) :) :)

OpenStudy (anonymous):

@creamsupreme plz help me

OpenStudy (anonymous):

i'm sorry i cant help you i haven't done limits yet

OpenStudy (anonymous):

@creamsupreme don't worry i'm also beginner :)

OpenStudy (anonymous):

did you understand the paper from @Eyad

OpenStudy (anonymous):

yah... i got everything.. i started doing limits jst nearabout 20 days before .. now i know alot :)

OpenStudy (anonymous):

so apply the same thing to your question

OpenStudy (anonymous):

yah.. but i want the detail of finding the limit of my question.... i know the question and answer but not the intermediate details :)

OpenStudy (anonymous):

oh sorry can't help you with that :P

OpenStudy (anonymous):

@creamsupreme .. no problem at all .. :) :).. i don't know my all fans are offline today what happened to them :) :)

OpenStudy (anonymous):

i know what you mean

OpenStudy (anonymous):

sry @jatinbansalhot ,i g2g .. anyway i'am online now ,so about your question the first step u will do when u find any limit question is subs. x with the number and if the result is any fraction its denominator is 0 so the limit will be Infinity ,In your question we will substitute x with 2 and its gonna be 2-3/2^2-4=-1/0 = infinity , u got it :)

OpenStudy (anonymous):

@Eyad you are so nice thanks for ur kind help to me... i had followed ur limit image paper and i learned so helpful step .... thanks a lot Eyad... :) :) :) :)